OpenStudy (anonymous):

Carla Lopez deposits $3000 a year into her retirement account. If these funds have an average earning of 9% over the 40 years until her retirement, what would be the value of her retirement account?

OpenStudy (helder_edwin):

if P=$3000 is your first deposit after one year you have \[ \large P+rP=(1+r)P \] where r is the interest rate. in your case r=9%=0.09 after the second year \[ \large (1+r)P+r(1+r)P=(1+r)^2P \] After n years the wealth in your retirement account is \[ \large W=(1+r)^nP \]

OpenStudy (anonymous):

can you explain this a little better i'm really confused

OpenStudy (anonymous):

is the equation (1+.09)^40x3000?

OpenStudy (helder_edwin):

if you save $100 and your bank pays you a 9% interest rate, it means that they are paying you 9c for every $. how much do u have after one year?

OpenStudy (helder_edwin):

yes that is what u have to compute

OpenStudy (anonymous):

I got 94228.26 is this correct?

OpenStudy (helder_edwin):

yes

OpenStudy (anonymous):

I was reading the answer online and it gave me $3000 x 337.890=1,013,670 which is why I am confused because I don't see where they get 337.890

OpenStudy (helder_edwin):

ok

OpenStudy (helder_edwin):

i did not read your question correctly!

OpenStudy (anonymous):

Do you know how they got this?

OpenStudy (helder_edwin):

it says there's a deposit of $3000 every year!

OpenStudy (helder_edwin):

give a few minutes to check a book on finance

OpenStudy (anonymous):

ok thank you so much

OpenStudy (helder_edwin):

ok here is what i got

OpenStudy (helder_edwin):

what i did before is good for the first deposit of $3000

OpenStudy (helder_edwin):